# LARP: Learner-Agnostic Robust Data Prefiltering

### v1

@@ 1 Introduction
The availability of large, public datasets has been crucial for the recent successes of statistical and machine learning methods. For example, public benchmarks such as ImageNET [13] and GLUE [59] have enabled the creation of numerous pre-trained and competitive models, readily available for adoption and fine-tuning by practitioners in any industry. Similarly, public healthcare datasets, e.g. the public data on the COVID-19 pandemic released by the WHO [^1] https://data.who.int/dashboards/covid19/data, can serve as a valuable reference to medical professionals for estimating statistics about common diseases and treatments.
Despite their attractiveness, public datasets often contain noisy, inaccurate or even maliciously manipulated data [8]. Since statistical methods are vulnerable to data contamination, such data issues can damage the accuracy of the downstream learning procedures applied on top of public data. For example, the Common Crawl web corpus, a widely used dataset in foundation model training, is known to contain toxic, biased, and factually incorrect content [23, 17]. Models trained on such data, including early versions of GPT-3 [7], have consequently been shown to produce toxic or biased outputs [56]. Another impact of public data inaccuracies manifested during the COVID-19 pandemic, as estimates of the disease mortality based on (hastily-collected) early public data on the virus spread were later reported to be underestimates [42].
These issues motivate the question of whether and how public datasets should be prefiltered by the data provider, so as to protect the accuracy of downstream statistical and learning procedures (hereafter referred to as learners, for brevity) applied on top of the dataset. Intuitively, principled data prefiltering can improve overall model downstream quality and also aligns with recent calls for transparency in dataset creation [22] and for data-centric strategies in developing robust and ethical AI systems [39]. However, such prefiltering poses a new technical challenge: that of designing data prefiltering algorithms that are learner-agnostic, in the sense of protecting the accuracy of a wide specification of learners simultaneously.
@@
Our theoretical results indicate that the learner-agnostic risk generally becomes larger with increased heterogeneity of the specified learner set, leading to a loss in utility compared to the alternative of prefiltering data for each learner/use-case individually (learner-specific prefiltering). We term the average utility reduction across the learner set as the price of learner-agnostic prefiltering and study it via extensive experiments on the Adult [3] and CIFAR-10 [36] datasets, with label noise and spurious correlations data corruptions. Finally, we model the trade-off between the price of LARP and the cost of repeated (learner-specific) prefiltering within a game-theoretic framework and showcase benefits of LARP for large datasets.
@@ 2 Related work
The problem of learning reliable estimates/models from contaminated data has long been studied in the fields of robust statistics and machine learning [29, 11]. Numerous works have studied various contamination models in the context of statistical estimation [28, 29, 33, 15, 32, 14] and supervised learning [33, 4, 16]. Some commonly studied type of data corruption in machine learning include label noise [44, 48, 26, 45, 62] and spurious correlations [24, 53, 43, 51], which we consider in our experiments.
@@
Recent advancements in data-centric machine learning emphasize the critical role of data quality in model performance. Surveys by [61] and [60] provide comprehensive overviews of methodologies aimed at improving data quality through validation, cleaning, and maintenance. One line of work studies learner-agnostic methods in the context of data valuation [31, 34]. Another approach, which is focused on mitigating label noise, is the data pruning strategy [46]. While we are also interested in prefiltering procedures that improve data quality, we study a new optimization objective for this task: that of minimizing the maximum loss of set of pre-specified learning algorithms.
@@
A large body of literature studies methods for data quality control and moderation for ML data obtained via crowdsourcing [38, 2, 58, 54]. These works study specific models of data contamination, e.g. label noise, and develop techniques for data filtering with provable guarantees on the resulting data quality. Unlike these works, we only adopt limited assumptions on the type of corruption and seek to certify the quality of downstream learners directly.
Another line of work studies the impact of data moderation in the context of human learning and disinformation [25, 19, 30]. The social implications in such settings lead to many orthogonal considerations for moderators, in particular designing appropriate models of human learning, reducing polarization and diversity of the provided information. In contrast, we focus protecting the accuracy of downstream learning algorithms despite corrupted data.
@@
Our work is motivated by the increasing availability of open-access datasets for training, which necessitates their preprocessing and preparation for public use. This is an example of the increasing fragmentation of modern learning pipelines, as multiple stakeholders become involved in one or more stages of the ML pipeline, including data gathering, preprocessing, training and deployment. In particular, several works have focused on interactions between data providers and model creators in ML. Data delegation and valuation have been studied by [9, 52, 1] with the goal of developing appropriate contract theory and fair remuneration methods for outsourcing data-related tasks. Other works consider strategic interactions between foundation model creators and entities performing fine-tuning [37]. Our work studies a concrete form of pipeline fragmentation, in which data preprocessing is performed independently of model training. We adopt the perspective of the data moderator, tasked with filtering corrupted data to ensure downstream learner robustness.
@@ 3 Framework
$R_{agn}(F)$ :$\vcentcolon=\max_{l\in\mathcal{L}}R_{l}(F)=\min_{l\in\mathcal{L}}R_{l}(F)+\left(\max_{l\in\mathcal{L}}R_{l}(F)-\min_{l\in\mathcal{L}}R_{l}(F)\right).$ (2)
@@ 4.1 Setup
We assume the standard Huber contamination [28], in which the $\epsilon$-contaminated sample $S$ is drawn i.i.d. from a distribution $\mathcal{D}^{\prime}_{\theta}=(1-\epsilon)\mathcal{D}_{\theta}+\epsilon Q,$ where $Q\in\mathfrak{Q}$ is a noise distribution. Note that since $Q$ is unknown in practice any upper bound on $R_{agn}$ should hold for all possible $Q\in\mathfrak{Q}$.
@@
The set of learners $\mathcal{L}$ is a set of Huber estimators [28, 57, 50], which are parametrized by a scalar $\delta>0$. They are defined as $\hat{\theta}_{\delta}(S)\vcentcolon=\operatorname*{argmin}_{\hat{\theta}\in\mathbb{R}}\sum_{x\in S}H_{\delta}(x-\hat{\theta})$, where the Huber loss $H_{\delta}$ is defined as
@@ 4.3 Prefiltering mechanism based on quantiles
We now prove feasibility of LARP in the context of a scalar mean estimation for a Gaussian distribution $\mathcal{N}(\theta,\sigma^{2})$, by providing a prefiltering procedure with provable upper bounds on its learner-agnostic risk. This procedure is inspired by a popular informal rule for marking the tails of a sample as outliers [41]. This can be adapted to define the following outlyingness measure:
@@
The first term in Equation 4 is reminiscent of upper bounds in classic robust mean estimation (see e.g. [14]) and increases with the corruption rate $\epsilon$ and decreases with the sample size. The second term depends solely on the learner set and is always nonnegative.
We note that small values of $\delta\approx 0$ yield estimators close to the sample median, which is known to enjoy guarantees similar to the first term in Equation 4 [14]. Therefore, if all estimators in $\mathcal{L}$ use a small value of $\delta$ (i.e. all estimators are “good” for the considered learning problem), the bound is comparable to those in standard robust mean estimation. However, if some estimators use a large $\delta$, they are suboptimal for the problem and the bound is also larger as $\max_{\delta\in\Delta}\delta^{2}$ is large. This aligns with the arguments presented in Section 4.2, since any non-trivial bound on $R_{agn}$ must also take into account the set of downstream learners.
@@ 4.4 Empirical analysis of the Gaussian mean estimation setup
In this subsection we empirically explore the dependence of $R_{agn}$ on $\epsilon$ and $\Delta$ with $\mathcal{D}_{\theta}=\mathcal{N}(0,1)$, by analyzing three different prefiltering procedures. The first one is $F_{p}^{q}$ as defined in Section 4.3, the other two are analogously defined using z-score [41] and Stahel-Donoho outlyingness (SDO) [55, 18, 63]. Each procedure is being optimized over their own hyperparameter that controls the amount of data being removed. Full details of the experiments are described in Section B.1.
In Figure 1(a) we notice that the empirical trend exhibits superlinear growth, in alignment with the $\epsilon^{2}$ term in Theorem 4.2. This is also consistent with lower bounds that we find in [14]. We also observe that $F^{z}$ and $F^{sdo}$ outperform $F^{q}$ in this setup. In any case, our results show that learning is viable under the assumption of reasonable downstream learners. In Figure 1(b) we measure the second term in Equation 2 on a set =$\mathcal{L}=\left\{\delta_{1},\delta_{2}\right\}$ where $\delta_{1}=0.01$ and $\delta_{2}$ varies along the x-axis, thus creating varying levels of heterogeneity between the learners. We see that increasing learner heterogeneity leads to larger additional loss incurred by the models. Moreover, the behaviour of this term is different for the different prefiltering procedures, suggesting that the optimal choice might depend on the learner set. This aligns with the analysis presented in Section 4.2, where we argued that the design of procedures for LARP should depend on the downstream learner set. In all cases, we conclude that the effect of the learner set is statistically significant and the guarantees for downstream learners are inherently worse due to learner heterogeneity.
@@ 5.1 Price of learner-agnostic prefiltering on real-world data
In this section, we quantify the price of learner-agnostic prefiltering on datasets with inherent noise. Specifically, we conduct experiments in both the tabular and image classification settings, using two realistic noise models: i) Uniform Label Noise, which mimics scenarios where annotation is unreliable; and ii) Shortcuts [24], which simulates scenarios where training data contains spurious correlation features irrelevant to the labels, e.g. textual information embedded in the CT image patches. In all cases, the price $P$ is computed using the utility function $\mathcal{U}_{red}(R_{1},R_{2})\vcentcolon=(R_{1}-R_{2})/R_{2}$, that measures the relative increase in classification risk when learner-agnostic prefiltering is used.
@@
For image experiments we use CIFAR-10 [36], splitting its train set into train (80%) and validation (20%) sets and using the original test set. Tabular experiments utilize Adult [3], randomly split into train (64%), validation (16%), and test (20%) sets. The validation sets are utilized to select the hyperparameters (Adult), and perform early-stopping (CIFAR-10). In all experiments, we apply the noise to the train and validation sets, but evaluate on clean data to measure learner generalization.
@@
In label noise experiments, we use the model accuracy as our risk metric $R$, $\epsilon=30\%$ uniform label noise on all classes, and prefilter up to $p\%$ of the train and validation sets. Motivated by the practical observation that complex models tend to learn useful features first before overfitting to noisy data, the prefiltered samples were selected by maximizing the loss of chosen overparametrized models (ResNet-9 [27] on CIFAR-10, and 2 layer fully connected NN on Adult) trained on the combined train and evaluation sets with early-stopping. We found this procedure effective in practice, resulting in $>75\%$ of corrupted samples being filtered at the default value of $p=25\%$.
@@
For tabular data, we inject noise by maximizing the integer feature “education-num” in $85\%$ of the data for the class corresponding to the label “$>$50K”, leading to $\epsilon\approx 21\%$ contamination rate due to class imbalance. For image data, we set the color of a $3\times 3$ patch in the lower right corner of $85\%$ of the data in classes 0,2,4 according to their label, resulting in $\epsilon=25.5\%$ contamination rate. Based on the observation that models learn shortcut data quickly, in both cases we prefilter the top-$p$ percent of lowest-loss datapoints based on CNN (CIFAR-10) and Random Forests (Adult) models from the respective learner sets, but trained on the combined train and validation sets. We verify that this results in $>85\%$ of prefiltered shortcut data in practice at the default $p=25\%$. We use the macro-F1 score [40] as our risk metric $R$ due to the class imbalance of the prefiltered datasets.
@@
We evaluate LARP on sets of models with different noise robustness levels, which can intuitively be seen as the worst-case heterogeneity, as models with high robustness will achieve optimal risk under little-to-no prefiltering, while models with low robustness will achieve optimal risk under heavy prefiltering. As we train very large number of models, for CIFAR-10 we opt to use convolutional networks (CNNs) with an architecture (See Section B.2) optimized for fast training that still maintains good accuracy. As we observe that L2 regularization increases the robustness to noisy labels in practice, but is also susceptible to underfitting, in the label noise setting, we generate our set of models $\mathcal{L}$ by varying the CNNs’ L2 regularization parameter in the range $[3e{-3},2e{-2}]$. For shortcut experiments we regularize the effect of the shortcuts on the training explicitly by reducing the gradient contribution of the shortcut patch pixels by a factor in the range $[1e{-6},1]$. For Adult experiments, we opt to use learners based on diverse set of machine learning model types — XGBoost [10], AdaBoost [20], LogitBoost [21], Bagging [5], RandomForest [6], SVM [12], as well as neural networks with two hidden layers and ReLU activations; in order to show that our conclusions generalize beyond families of models differing only by a single continuous parameter. For each model type we select two sets of hyperparameters — robust and brittle, resulting in 14 total models in $\mathcal{L}$. The robust ones were selected by minimizing the noisy validation data risk on 20 random hyperparameter configurations, while the brittle models were optimized to increase the difference between the risks on the prefiltered validation sets of two versions of the model – one trained on prefiltered train data, and one trained on noisy train data. Our brittle models intuitively represent learners that chose their parameters for maximal accuracy, disregarding robustness completely.
We train all neural networks using PyTorch [47] with Adam optimizer [35] and the remaining models are trained using Scikit-learn [49]. In all shortcut experiments learners reweight their losses to account for class imbalance of the prefiltered sets. Full experimental details can be found in Section B.2.

### v3

@@ 1 Introduction
The availability of large, public datasets has been crucial for the recent successes of statistical and machine learning methods. For example, public benchmarks such as ImageNET (Deng et al., 2009) and GLUE (Wang et al., 2019) have enabled the creation of numerous pre-trained and competitive models, readily available for adoption and fine-tuning by practitioners in any industry. Similarly, public healthcare datasets, e.g. the public data on the COVID-19 pandemic released by the WHO [^1] https://data.who.int/dashboards/covid19/data, can serve as a valuable reference to medical professionals for estimating statistics about common diseases and treatments.
Despite their attractiveness, public datasets often contain noisy, inaccurate or even maliciously manipulated data (Carlini et al., 2024). Since statistical methods are vulnerable to data contamination, such data issues can damage the accuracy of the downstream learning procedures applied on top of public data. For example, the Common Crawl web corpus, a widely used dataset in foundation model training, is known to contain toxic, biased, and factually incorrect content (Gehman et al., 2020; Dodge et al., 2021). Models trained on such data, including early versions of GPT-3 (Brown et al., 2020), have consequently been shown to produce toxic or biased outputs (Stiennon et al., 2020). Another impact of public data inaccuracies manifested during the COVID-19 pandemic, as estimates of the disease mortality based on (hastily-collected) early public data on the virus spread were later reported to be underestimates (Msemburi et al., 2023).
These issues motivate the question of whether and how public datasets should be prefiltered by the data provider, so as to protect the accuracy of downstream statistical and learning procedures (hereafter referred to as learners, for brevity) applied on top of the dataset. Intuitively, principled data prefiltering can improve overall model downstream quality and also aligns with recent calls for transparency in dataset creation (Gebru et al., 2021) and for data-centric strategies in developing robust and ethical AI systems (Liang et al., 2022). However, such prefiltering poses a new technical challenge: that of designing data prefiltering algorithms that are learner-agnostic, in the sense of protecting the accuracy of a wide specification of learners simultaneously.
@@
Our theoretical results indicate that the learner-agnostic risk generally becomes larger with increased heterogeneity of the specified learner set, leading to a loss in utility compared to the alternative of prefiltering data for each learner/use-case individually (learner-specific prefiltering). We term the average utility reduction across the learner set as the price of learner-agnostic prefiltering and study it via extensive experiments on the Adult (Becker & Kohavi, 1996) and CIFAR-10 (Krizhevsky, 2009) datasets, with label noise and spurious correlations data corruptions. Finally, we model the trade-off between the price of LARP and the cost of repeated (learner-specific) prefiltering within a game-theoretic framework and showcase benefits of LARP for large datasets.
@@ 2 Related work
The problem of learning reliable estimates/models from contaminated data has long been studied in the fields of robust statistics and machine learning (Huber, 2004; Cinà et al., 2023). Numerous works have studied various contamination models in the context of statistical estimation (Huber, 1964, 2004; Kearns & Li, 1988; Diakonikolas et al., 2019a; Kane et al., 2024; Diakonikolas & Kane, 2023) and supervised learning (Kearns & Li, 1988; Biggio et al., 2012; Diakonikolas et al., 2019b). Some commonly studied type of data corruption in machine learning include label noise (Natarajan et al., 2013; Patrini et al., 2017; Han et al., 2018; Northcutt et al., 2021; Zhang & Sabuncu, 2018) and spurious correlations (Geirhos et al., 2020; Shah et al., 2020; Nam et al., 2020; Sagawa et al., 2020), which we consider in our experiments.
@@
Recent advancements in data-centric machine learning emphasize the critical role of data quality in model performance. Surveys by (Zha et al., 2025) and (Whang et al., 2023) provide comprehensive overviews of methodologies aimed at improving data quality through validation, cleaning, and maintenance. One line of work studies learner-agnostic methods in the context of data valuation (Just et al., 2023; Kessler et al., 2025). Another approach, which is focused on mitigating label noise, is the data pruning strategy (Park et al., 2023). While we are also interested in prefiltering procedures that improve data quality, we study a new optimization objective for this task: that of minimizing the maximum loss of set of pre-specified learning algorithms.
@@
A large body of literature studies methods for data quality control and moderation for ML data obtained via crowdsourcing (Lease, 2011; Awasthi et al., 2017; Vaughan, 2018; Sheng & Zhang, 2019). These works study specific models of data contamination, e.g. label noise, and develop techniques for data filtering with provable guarantees on the resulting data quality. Unlike these works, we only adopt limited assumptions on the type of corruption and seek to certify the quality of downstream learners directly.
Another line of work studies the impact of data moderation in the context of human learning and disinformation (Haghtalab et al., 2021; Dwork et al., 2024; Huleihel & Refael, 2024). The social implications in such settings lead to many orthogonal considerations for moderators, in particular designing appropriate models of human learning, reducing polarization and diversity of the provided information. In contrast, we focus protecting the accuracy of downstream learning algorithms despite corrupted data.
@@
Our work is motivated by the increasing availability of open-access datasets for training, which necessitates their preprocessing and preparation for public use. This is an example of the increasing fragmentation of modern learning pipelines, as multiple stakeholders become involved in one or more stages of the ML pipeline, including data gathering, preprocessing, training and deployment. In particular, several works have focused on interactions between data providers and model creators in ML. Data delegation and valuation have been studied by (Chen et al., 2022; Saig et al., 2023; Ananthakrishnan et al., 2024) with the goal of developing appropriate contract theory and fair remuneration methods for outsourcing data-related tasks. Other works consider strategic interactions between foundation model creators and entities performing fine-tuning (Laufer et al., 2024). Our work studies a concrete form of pipeline fragmentation, in which data preprocessing is performed independently of model training. We adopt the perspective of the data moderator, tasked with filtering corrupted data to ensure downstream learner robustness.
@@ 3 Framework
$\begin{split}R_{agn}(F)&\vcentcolon=\max_{l\in\mathcal{L}}R_{l}(F)=\\ &=\min_{l\in\mathcal{L}}R_{l}(F)+\left(\max_{l\in\mathcal{L}}R_{l}(F)-\min_{l\in\mathcal{L}}R_{l}(F)\right).\end{split}$ (2)
@@ 4.1 Setup
We assume the standard Huber contamination (Huber, 1964), in which the $\epsilon$-contaminated sample $S$ is drawn i.i.d. from a distribution $\mathcal{D}^{\prime}_{\theta}=(1-\epsilon)\mathcal{D}_{\theta}+\epsilon Q,$ where $Q\in\mathfrak{Q}$ is a noise distribution. Note that since $Q$ is unknown in practice any upper bound on $R_{agn}$ should hold for all possible $Q\in\mathfrak{Q}$.
@@
The set of learners $\mathcal{L}$ is a set of Huber estimators (Huber, 1964; Sun et al., 2020; Pensia et al., 2024), which are parametrized by a scalar $\delta>0$. They are defined as $\hat{\theta}_{\delta}(S)\vcentcolon=\operatorname*{argmin}_{\hat{\theta}\in\mathbb{R}}\sum_{x\in S}H_{\delta}(x-\hat{\theta})$, where the Huber loss $H_{\delta}$ is defined as
@@ 4.3 Prefiltering mechanism based on quantiles
We now prove feasibility of LARP in the context of a scalar mean estimation for a Gaussian distribution $\mathcal{N}(\theta,\sigma^{2})$, by providing a prefiltering procedure with provable upper bounds on its learner-agnostic risk. This procedure is inspired by a popular informal rule for marking the tails of a sample as outliers (Maronna et al., 2006). This can be adapted to define the following outlyingness measure:
@@
The first term in Equation 5 is reminiscent of upper bounds in classic robust mean estimation (see e.g. (Diakonikolas & Kane, 2023)) and increases with the corruption rate $\epsilon$ and decreases with the sample size. The second term depends solely on the learner set and is always nonnegative.
We note that small values of $\delta\approx 0$ yield estimators close to the sample median, which is known to enjoy guarantees similar to the first term in Equation 5 (Diakonikolas & Kane, 2023). Therefore, if all estimators in $\mathcal{L}$ use a small value of $\delta$ (i.e. all estimators are “good” for the considered learning problem), the bound is comparable to those in standard robust mean estimation. However, if some estimators use a large $\delta$, they are suboptimal for the problem and the bound is also larger as $\max_{\delta\in\Delta}\delta^{2}$ is large. This aligns with the arguments presented in Section 4.2, since any non-trivial bound on $R_{agn}$ must also take into account the set of downstream learners.
@@ 4.4 Empirical analysis of the Gaussian mean estimation setup
In this subsection we empirically explore the dependence of $R_{agn}$ on $\epsilon$ and $\Delta$ with $\mathcal{D}_{\theta}=\mathcal{N}(0,1)$, by analyzing three different prefiltering procedures. The first one is $F_{p}^{q}$ as defined in Section 4.3, the other two are analogously defined using z-score (Maronna et al., 2006) and Stahel-Donoho outlyingness (SDO) (Stahel, 1981; Donoho, 1982; Zuo et al., 2004). Each procedure is being optimized over their own hyperparameter that controls the amount of data being removed. Full details of the experiments are described in Section B.1.
In Figure 1(a) we notice that the empirical trend exhibits superlinear growth, in alignment with the $\epsilon^{2}$ term in Theorem 4.2. This is also consistent with lower bounds that we find in (Diakonikolas & Kane, 2023). We also observe that $F^{z}$ and $F^{sdo}$ outperform $F^{q}$ in this setup. In any case, our results show that learning is viable under the assumption of reasonable downstream learners. In Figure 1(b) we measure the second term in Equation 2 on a set =$\mathcal{L}=\left\{\delta_{1},\delta_{2}\right\}$ where $\delta_{1}=0.01$ and $\delta_{2}$ varies along the x-axis, thus creating varying levels of heterogeneity between the learners. We see that increasing learner heterogeneity leads to larger additional loss incurred by the models. Moreover, the behaviour of this term is different for the different prefiltering procedures, suggesting that the optimal choice might depend on the learner set. This aligns with the analysis presented in Section 4.2, where we argued that the design of procedures for LARP should depend on the downstream learner set. In all cases, we conclude that the effect of the learner set is statistically significant and the guarantees for downstream learners are inherently worse due to learner heterogeneity.
@@ 5.1 Price of learner-agnostic prefiltering on real-world data
In this section, we quantify the price of learner-agnostic prefiltering on datasets with inherent noise. Specifically, we conduct experiments in both the tabular and image classification settings, using two realistic noise models: i) Uniform Label Noise, which mimics scenarios where annotation is unreliable; and ii) Shortcuts (Geirhos et al., 2020), which simulates scenarios where training data contains spurious correlation features irrelevant to the labels, e.g. textual information embedded in the CT image patches. In all cases, the price $P$ is computed using the utility function $\mathcal{U}_{red}(R_{1},R_{2})\vcentcolon=(R_{1}-R_{2})/R_{2}$, that measures the relative increase in classification risk when learner-agnostic prefiltering is used.
@@
For image experiments we use CIFAR-10 (Krizhevsky, 2009), splitting its train set into train (80%) and validation (20%) sets and using the original test set. Tabular experiments utilize Adult (Becker & Kohavi, 1996), randomly split into train (64%), validation (16%), and test (20%) sets. The validation sets are utilized to select the hyperparameters (Adult), and perform early-stopping (CIFAR-10). In all experiments, we apply the noise to the train and validation sets, but evaluate on clean data to measure learner generalization.
@@
In label noise experiments, we use the model accuracy as our risk metric $R$, $\epsilon=30\%$ uniform label noise on all classes, and prefilter up to $p\%$ of the train and validation sets. Motivated by the practical observation that complex models tend to learn useful features first before overfitting to noisy data, the prefiltered samples were selected by maximizing the loss of chosen overparametrized models (ResNet-9 (He et al., 2016) on CIFAR-10, and 2 layer fully connected NN on Adult) trained on the combined train and evaluation sets with early-stopping. We found this procedure effective in practice, resulting in $>75\%$ of corrupted samples being filtered at the default value of $p=25\%$.
@@
For tabular data, we inject noise by maximizing the integer feature “education-num” in $85\%$ of the data for the class corresponding to the label “$>$50K”, leading to $\epsilon\approx 21\%$ contamination rate due to class imbalance. For image data, we set the color of a $3\times 3$ patch in the lower right corner of $85\%$ of the data in classes 0,2,4 according to their label, resulting in $\epsilon=25.5\%$ contamination rate. Based on the observation that models learn shortcut data quickly, in both cases we prefilter the top-$p$ percent of lowest-loss datapoints based on CNN (CIFAR-10) and Random Forests (Adult) models from the respective learner sets, but trained on the combined train and validation sets. We verify that this results in $>85\%$ of prefiltered shortcut data in practice at the default $p=25\%$. We use the macro-F1 score (Manning, 2009) as our risk metric $R$ due to the class imbalance of the prefiltered datasets.
@@
We evaluate LARP on sets of models with different noise robustness levels, which can intuitively be seen as the worst-case heterogeneity, as models with high robustness will achieve optimal risk under little-to-no prefiltering, while models with low robustness will achieve optimal risk under heavy prefiltering. As we train very large number of models, for CIFAR-10 we opt to use convolutional networks (CNNs) with an architecture (See Section B.2) optimized for fast training that still maintains good accuracy. As we observe that L2 regularization increases the robustness to noisy labels in practice, but is also susceptible to underfitting, in the label noise setting, we generate our set of models $\mathcal{L}$ by varying the CNNs’ L2 regularization parameter in the range $[3e{-3},2e{-2}]$. For shortcut experiments we regularize the effect of the shortcuts on the training explicitly by reducing the gradient contribution of the shortcut patch pixels by a factor in the range $[1e{-6},1]$. For Adult experiments, we opt to use learners based on diverse set of machine learning model types — XGBoost (Chen & Guestrin, 2016), AdaBoost (Freund & Schapire, 1997), LogitBoost (Friedman et al., 2000), Bagging (Breiman, 1996), RandomForest (Breiman, 2001), SVM (Cortes & Vapnik, 1995), as well as neural networks with two hidden layers and ReLU activations; in order to show that our conclusions generalize beyond families of models differing only by a single continuous parameter. For each model type we select two sets of hyperparameters — robust and brittle, resulting in 14 total models in $\mathcal{L}$. The robust ones were selected by minimizing the noisy validation data risk on 20 random hyperparameter configurations, while the brittle models were optimized to increase the difference between the risks on the prefiltered validation sets of two versions of the model – one trained on prefiltered train data, and one trained on noisy train data. Our brittle models intuitively represent learners that chose their parameters for maximal accuracy, disregarding robustness completely.
We train all neural networks using PyTorch (Paszke et al., 2019) with Adam optimizer (Kingma & Ba, 2015) and the remaining models are trained using Scikit-learn (Pedregosa et al., 2011). In all shortcut experiments learners reweight their losses to account for class imbalance of the prefiltered sets. Full experimental details can be found in Section B.2.